OpenStudy (anonymous):

[(x^2+2x) - 2y + 1 ] NEED HELP PLEASE

OpenStudy (anonymous):

what are you required to do?

OpenStudy (anonymous):

factorise? simplify? expand?

OpenStudy (anonymous):

@mo7892 simplify. :)

OpenStudy (anonymous):

ok am assuming the y has to be on one side and x on the other side

OpenStudy (anonymous):

then? @mo7892

OpenStudy (anonymous):

[(x²+2x) - 2y + 1 ]=0 is that how the equation is written in the book or paper

OpenStudy (anonymous):

[(x²+2x) - 2y + 1 ]=0 (x²+2x) +1 =2y 1/2 (x²+2x) +1/2 =y

OpenStudy (anonymous):

do you get it so far

OpenStudy (anonymous):

x²/2+x+1/2=y from here you can factorise but i don't think they have asked for that.